# A characterization of zero entropy loosely Bernoulli flows via FK-pseudometric

### v1

@@ 3.2. Feldman-Katok pseudometric for flows
Inspired by the $\bar{f}$-metric Kwietniak and Łącka introduced in [KŁ17] the FK-pseudometric for discrete-time systems. Indpendently, Ratner [Rat17] proposed a notion of matching between segments of orbits for flows that can be seen as a continuous-time version of the $\bar{f}_{n}$-metric with respect to partitions (see section 4 for details). Motivated by these works we propose a definition of the Feldman-Katok pseudometric for flows.
@@
Let $x,y\in X$, $t>1$, $\varepsilon>0$ and $\delta>0$. We say that $x$ and $y$ are $\widetilde{(t,\varepsilon,\delta)}$-matchable if there exist measurable sets $A,A^{\prime}\subseteq[0,t]$ with $\lambda(A)>(1-\varepsilon)t$, $\lambda(A^{\prime})>(1-\varepsilon)t$ and an increasing absolutely continuous onto map $h:A\to A^{\prime}$ satisfying for all $s\in A$:
@@
Let $E\subset X$ be a Borel set. We will show that $\mu_{x,t}(E)\leq\mu_{y,t}(E^{\beta})+\beta$ where $\beta=\max\{\delta,2\varepsilon\}$. For $z\in X$ and $B\subset X$ we denote $B(z,t)=\{s\in[0,t]:\varphi^{s}(z)\in B\}$.
Let $h:A\to A^{\prime}$ a $\widetilde{(t,\varepsilon,\delta)}$-matching between $x$ and $y$. Since $\lambda(A)>(1-\varepsilon)t$ we have
@@
If $s\in A\cap E(x,t)$, then $\varphi^{h(s)}(y)\in E^{\delta}$, which means $h(s)\in E^{\delta}(y,t)$. Thus $h(E(x,t)\cap A)\subset E^{\delta}(y,t)$ and consequently, putting it all together we obtain:
@@
The $\tilde{f}_{\delta}$-pseudodistance is defined as:
@@
The Feldman-Katok pseudometric (FK-pseudometric for short) is given by:
@@
In several cases, topological and ergodic properties of a flow cannot be projected onto all time-t maps. For example, this happens with such properties like minimality, ergodicity and unique ergodicity. On the other hand, sometimes it is enough that a property holds for a time-t map for some $t$ to guarantee that it holds for the flow. Examples of the later properties are ergodicity and unique ergodicity (see [FH19] for more details).
In both, Theorem 1 and Theorem 2, we focus on zero FK-distance between pairs of points. We will show that if the FK-distance between a pair of points with respect to a time-t map vanishes for some $t$, then the FK-distance between these points with respect to the flow also vanishes.
@@
Let $t_{0}>n_{0}$ be such that $\frac{t_{0}\varepsilon}{8}>1$. Fix $t\geq t_{0}$. Then, there exists $n\in\mathbb{N}$ with $n\geq n_{0}$ such that $n<t\leq n+1$. Since $\bar{f}_{n+1,\delta}(x,y)\leq\delta$, there exists $\pi\colon D\to D^{\prime}$ an $(n+1,\delta)$-matching between $x$ and $y$ with $|\pi|\geq(1-\delta)(n+1)$.
@@
$|D_{0}|$ $\geq(1-\delta)(n+1)-4$
$\geq(1-\delta)t-4$
$>(1-\varepsilon)t.$
@@
Let $h\colon A\to A^{\prime}$ be the isometric translation from $[k,k+1)$ to $[\pi(k),\pi(k)+1)$ for each $k\in D_{0}$.
By construction we have $A,A^{\prime}\subset[0,t]$ and $\lambda(A)=\lambda(A^{\prime})=|D_{0}|>(1-\varepsilon)t$. Since $d(\varphi^{k}(x),\varphi^{\pi(k)}(y))<\delta$ for every $k\in D_{0}$, from the choice of $\delta$ we conclude that $d(\varphi^{s}(x),\varphi^{h(s)}(y))<\varepsilon$ for every $s\in A$. Therefore, $h$ is a $\widetilde{(t,\varepsilon,\varepsilon)}$-matching between $x$ and $y$ which completes the proof. ∎
@@
Note that $d_{T}(\widehat{w},\widehat{z})=d(w,z)$ for every $w,z\in X$ and that $\varphi^{1}(\widehat{z})=\widehat{T(z)}$ for every $z\in X$.
Let $\operatorname{\rho^{1}_{FK}}$ denote the FK-pseudometric on $X_{T}$ with respect to $\varphi^{1}$. So $\operatorname{\rho_{FK}}(x,y)=0$ implies $\operatorname{\rho^{1}_{FK}}(\widehat{x},\widehat{y})=0$ and consequently, by Proposition 3.14, $\operatorname{\tilde{\rho}_{FK}}(\widehat{x},\widehat{y})=0$. ∎
@@ 4. Loosely Kronecker flows
To any ordering of a given partition ${\mathcal{P}}=\{P_{1},...,P_{n}\}$ we can associate a function on $X$ defined by ${\mathcal{P}}(x)=j$ for $x\in P_{j}$.
@@
where the infimum is taken over all possible orderings for the partitions ${\mathcal{P}}$ and ${\mathcal{Q}}$. In case $|{\mathcal{P}}|\neq|{\mathcal{Q}}|$ we add empty cells to the partition with less elements in order to have two partitions with the same cardinality. If we identify partitions ${\mathcal{P}},{\mathcal{Q}}$ such that $d_{\mu}({\mathcal{P}},{\mathcal{Q}})=0$, then $d_{\mu}$ is indeed a metric on $\operatorname{\mathbf{P}}(X)$ (here we abuse the terminology and treat equivalence classes of partitions as partitions).
@@
In [Rat17], the author proved that for ergodic flows the quantity $e(\Phi,u)$ remains the same if in the definition of matching we only demand that $h$ is an increasing measure-preserving map. It is also proved that this quantity is invariant under Kakutani equivalence. Then a characterization of loosely Kronecker flows using the invariant $e(\Phi,u)$ is the following:
@@
Since the FK-pseudometric is a topological notion, we focus on some partitions with good topological properties and we show that it is enough to study them when computing $e(\Phi,u)=0$.
We start fixing ${\mathcal{U}}$ a countable basis for the topology of $X$. We can assume without loss of generality that ${\mathcal{U}}$ is closed under finite unions and intersections. We say that a partition is ${\mathcal{U}}$-regular if at most one cell does not belong to ${\mathcal{U}}$.
@@
For a given $\mu\in{\mathcal{M}}(X)$, we say that a partition ${\mathcal{P}}$ is $\varepsilon$-essentially open if it is ${\mathcal{U}}$-regular and if there is a cell $P\in{\mathcal{P}}$ that does not belong to ${\mathcal{U}}$, then it satisfies $\mu(P)<\varepsilon$.
Besides having good topological properties, for each $\mu\in{\mathcal{M}}(X)$ the family of $\varepsilon$-essentially partitions constitutes a dense set in the space of measurable partitions.
@@
Let ${\mathcal{P}}=\{P_{1},...,P_{n}\}\in\operatorname{\mathbf{P}}(X)$. Fix $0<\delta<\varepsilon$. By regularity of $\mu$, for each $i\in\{1,...,n\}$ there exists a compact $K_{i}\subset P_{i}$ such that $\mu(P_{i}\backslash K_{i})<\frac{\delta}{4n}$. Take $\{Q_{1},...,Q_{n}\}\subset{\mathcal{U}}$ disjoint such that $K_{i}\subset Q_{i}$ and $\mu(Q_{i}\backslash K_{i})<\frac{\delta}{4n}$ for $i\in\{1,...,n\}$. Set $Q_{n+1}=X\backslash\cup_{i=1}^{n}Q_{i}$ and ${\mathcal{Q}}=\{Q_{1},...,Q_{n+1}\}$.
Observe that since $Q_{n+1}\subset\cup_{i=1}^{n}P_{i}\backslash K_{i}$ we have $\mu(Q_{n+1})<\frac{\delta}{4}<\varepsilon$ which means that ${\mathcal{Q}}$ is $\varepsilon$-essentially open. Moreover, $\mu(P_{j}\triangle Q_{j})<\frac{\delta}{2n}$ for every $j\in\{1,...,n\}$ and therefore $d_{\mu}({\mathcal{P}},{\mathcal{Q}})<\delta$ which completes the proof. ∎
One direction of Theorem 1 is a direct consequence of García-Ramos and Kwietniak’s result (Theorem 4.5 of [GRK22]) together with some properties of the FK-pseudometric already presented.
@@
Assume that ${\mathcal{P}}$ and ${\mathcal{Q}}$ are ordered optimally with respect to $d_{\mu}$. Let $D=\{z\in X:{\mathcal{P}}(z)\neq{\mathcal{Q}}(z)\}$. For any $z\in X$ and $t\in\mathbb{R}^{+}$ we will denote by $E(z,t,D)=\{s\in[0,t]:\varphi^{s}(z)\notin D\}$. Since $\mu$ is ergodic there exists $H\subset\operatorname{Gen}(\mu)$ and $t_{0}>0$ such that $\mu(H)>1-\delta$ and for every $z\in H$ and $t\geq t_{0}$ we have:
(4.1) $|\mu_{z,t}(D)-d_{\mu}({\mathcal{P}},{\mathcal{Q}})|=|\mu_{z,t}(D)-\mu(D)|<\frac{\delta}{2}.$
@@
(4.2) $\frac{\lambda(E(x,t,D))}{t}>1-d_{\mu}({\mathcal{P}},{\mathcal{Q}})-\frac{\delta}{2},$
and the same holds for $\lambda(E(y,t,D))$. Using these estimates together with the control of the measure by the derivative of $h$ (see condition (1) in Definition 4.1) and the fact that $\lambda(A_{0}^{\prime})>1-\varepsilon$ we have
(4.3) $\frac{\lambda(h^{-1}(A_{0}^{\prime}\cap E(y,t,D)))}{t}\geq\frac{\lambda(A_{0}^{\prime}\cap E(y,t,D))}{t}(1-\varepsilon)>1-2\varepsilon-d_{\mu}({\mathcal{P}},{\mathcal{Q}})-\frac{\delta}{2}.$
@@
The last ingredient is due to Ornstein, Rudolph and Weiss (see Chapter 7 of [ORW82]). The result was initially obtained by Rudolph in [Rud78] for loosely Bernoulli flows (the positive entropy case) and a modification of the argument led to an analogous result for loosely Kronecker flows.
@@
Let $\Phi$ be a continuous flow on $X$ and $\mu\in\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$. The measure-preserving flow $(X,\Phi,\mu)$ is loosely Kronecker if and only if there exists a Borel set $H\subset X$ such that $\mu(H)=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
@@
By the discrete-time result analogous to Theorem 1 (Theorem 4.5 of [GRK22]) we obtain a Borel set $H\subset X$ with $\mu(H)=1$ such that $\operatorname{\rho_{FK}}(x,y)=0$ with respect to $\varphi^{t}$ for every $x,y\in H$ By Proposition 3.14 we obtain $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
For the converse we fix $\varepsilon>0$, a partition ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ and a Borel set $H_{0}\subset X$ such that $\mu(H_{0})=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H_{0}$.
From the density of essentially open partitions there exists an $\varepsilon$-essentially open partition ${\mathcal{Q}}=\{Q_{0},...,Q_{n}\}$ such that $d_{\mu}({\mathcal{P}},{\mathcal{Q}})<\varepsilon$. We assume without loss of generality that $Q_{0}$ is the compact cell and take an open set $V_{0}$ containing $Q_{0}$ such that $\mu(V_{0})<\varepsilon$.
Let $0<\delta<\varepsilon$ be a Lebesgue number for the open cover ${\mathcal{W}}=\{V_{0},Q_{1},...,Q_{n}\}$ of $X$.
Since $\mu$ is ergodic, there exist a compact set $H_{1}\subset\operatorname{Gen}(\mu)\cap H_{0}$ and $t_{1}>1$ such that $\mu(H_{1})>1-\delta$ and $\mu_{x,t}(V_{0})<\varepsilon$ for every $t\geq t_{1}$ and $x\in H_{1}$.
Fix $x_{0}\in H_{1}$. We observe that $\tilde{f}_{t,\delta}(x_{0},\cdot)$ is continuous for every $t\geq 1$, so $\tilde{f}_{\delta}(x_{0},\cdot)$ is measurable. Thus there exist a compact set $H_{2}\subset H_{1}$ and $t_{2}\geq t_{1}$ such that $\mu(H_{2})>1-\delta$ and $\tilde{f}_{t,\delta}(x_{0},y)<\delta$ for every $t\geq t_{2}$ and $y\in H_{2}$.
For $t\geq t_{2}$ and $y\in H_{2}$ let $h\colon A_{0}\to A_{0}^{\prime}$ be a $\widetilde{(t,\delta,\delta)}$-matching between $x_{0}$ and $y$. Set $A\subset A_{0}$ as:
@@
Using an argument similar to the one in the in the proof of Lemma 4.5 together with the fact that $\delta<\varepsilon$ we can see that $h|_{A}\colon A\to h(A)$ is a $(t,3\varepsilon,{\mathcal{Q}})$-matching between $x_{0}$ and $y$.
Lastly, we apply Lemma 4.5 and assume without loss of generality that the obtained set $H$ is contained in $H_{2}$ and $t_{0}\geq t_{2}$. So for every $y\in H$ and $t\geq t_{0}$ we have
@@
This means that $H\subset B_{t}(x_{0},15\varepsilon,{\mathcal{P}})$, and consequently, $K_{t}(15\varepsilon,{\mathcal{P}})=1$ for every $t\geq t_{0}$. Since ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ and $\varepsilon>0$ were arbitrary, we have $e(\Phi,u)=0$ for every $u\in U$. Using Theorem 4.2 we conclude that $(X,\Phi,\mu)$ is loosely Kronecker. ∎

### v2

@@ 3.2. Feldman-Katok pseudometric for flows
Inspired by the $\bar{f}$-metric Kwietniak and Łącka introduced in [KŁ17] the FK-pseudometric for discrete-time systems. Independently, Ratner [Rat17] proposed a notion of matching between segments of orbits for flows that can be seen as a continuous-time version of the $\bar{f}_{n}$-metric with respect to partitions (see section 4 for details). Motivated by these works we propose a definition of the Feldman-Katok pseudometric for flows.
@@
Let $x,y\in X$, $t>1$, $10<\varepsilon<$ 1 and $\delta>0$. We say that $x$ and $y$ are $\widetilde{(t,\varepsilon,\delta)}$-matchable if there exist measurable sets $A,A^{\prime}\subseteq[0,t]$ with $\lambda(A)>(1-\varepsilon)t$, $\lambda(A^{\prime})>(1-\varepsilon)t$ and an increasing absolutely continuous onto map $h:A\to A^{\prime}$ satisfying for all $s\in A$:
@@
Let $E\subseteq X$ be a Borel set. We will show that $\mu_{x,t}(E)\leq\mu_{y,t}(E^{\beta})+\beta$ where $\beta=\max\{\delta,2\varepsilon\}$. For $z\in X$ and $B\subseteq X$ we denote $B(z,t)=\{s\in[0,t]:\varphi^{s}(z)\in B\}$.
Let $h:A\to A^{\prime}$ be a $\widetilde{(t,\varepsilon,\delta)}$-matching between $x$ and $y$. Since $\lambda(A)>(1-\varepsilon)t$ we have
@@
If $s\in A\cap E(x,t)$, then $\varphi^{h(s)}(y)\in E^{\delta}$, which means $h(s)\in E^{\delta}(y,t)$. Thus $h(E(x,t)\cap A)\subseteq E^{\delta}(y,t)$ and consequently, putting it all together we obtain
@@
The $\tilde{f}_{\delta}$-pseudodistance is defined as
@@
The Feldman-Katok pseudometric (FK-pseudometric for short) is given by
@@
In general, we can only guarantee that topological and ergodic properties of a flow are projected onto many but not all time-$t$ maps. This happens with properties like minimality, ergodicity and unique ergodicity.
Proposition 3.14 (Proposition 4 in [Fay00]).
Let $\Phi$ be a continuous flow. If $\Phi$ is transitive (resp. minimal), then $\varphi^{t}$ is transitive (resp. minimal) for a dense $G_{\delta}$ set of $t\in\mathbb{R}$.
A similar result in the ergodic setting was proved by Pugh and Shub.
Proposition 3.15 (Theorem 1 in [PS71]).
Let $\Phi$ be a continuous flow on $X$ and $\mu\in\operatorname{{\mathcal{M}}_{\Phi}}(X)$. If $(X,\Phi,\mu)$ is ergodic, then $(X,\varphi^{t},\mu)$ is ergodic for all but countably many $t\in\mathbb{R}$.
The proof (due to Veech) of Theorem 3.3.34 in [FH19] implies the next result.
Proposition 3.16.
If a continuous flow $\Phi$ is uniquely ergodic, then $\varphi^{t}$ is uniquely ergodic for all but countably many $t\in\mathbb{R}$.
On the other hand, sometimes it is enough that a property holds for a time-$t$ map for some $t$ to guarantee that it holds for the flow. Examples of the latter properties are ergodicity and unique ergodicity (see [FH19] for more details).
In both Theorem 1 and Theorem 2, we focus on zero FK-distance between pairs of points. We will show that if the FK-distance between a pair of points with respect to a time-$t$ map vanishes for some $t$, then the FK-distance between these points with respect to the flow also vanishes.
@@
Let $t_{0}>n_{0}$ be such that $\frac{t_{0}\varepsilon}{4}>1$. Fix $t\geq t_{0}$. Then, there exists $n\in\mathbb{N}$ with $n\geq n_{0}$ such that $n<t\leq n+1$. Since $\bar{f}_{n+1,\delta}(x,y)\leq\delta$, there exists $\pi\colon D\to D^{\prime}$ an $(n+1,\delta)$-matching between $x$ and $y$ with $|\pi|\geq(1-\delta)(n+1)$.
@@
$|D_{0}|\geq(1-\delta)(n+1)-2\geq(1-\delta)t-2>(1-\varepsilon)t.$
@@
Let $h\colon A\to A^{\prime}$ be the isometric translation from $[k,k+1)$ to $[\pi(k),\pi(k)+1)$ for each $k\in D_{0}$. By construction we have $A,A^{\prime}\subseteq[0,t]$ and $\lambda(A)=\lambda(A^{\prime})=|D_{0}|>(1-\varepsilon)t$. Since $d(\varphi^{k}(x),\varphi^{\pi(k)}(y))<\delta$ for every $k\in D_{0}$, from the choice of $\delta$ we conclude that $d(\varphi^{s}(x),\varphi^{h(s)}(y))<\varepsilon$ for every $s\in A$. Therefore, $h$ is a $\widetilde{(t,\varepsilon,\varepsilon)}$-matching between $x$ and $y$ which completes the proof. ∎
@@
Let $\operatorname{\rho^{1}_{FK}}$ denote the FK-pseudometric on $X_{T}$ with respect to $\varphi^{1}$. Note that $d_{T}(\widehat{w},\widehat{z})=d(w,z)$ for every $w,z\in X$ and that $\varphi^{1}(\widehat{z})=\widehat{T(z)}$ for every $z\in X$. So $\operatorname{\rho_{FK}}(x,y)=0$ implies $\operatorname{\rho^{1}_{FK}}(\widehat{x},\widehat{y})=0$ and consequently, by Proposition 3.17, $\operatorname{\tilde{\rho}_{FK}}(\widehat{x},\widehat{y})=0$. ∎
@@ 4. Loosely Kronecker flows
To any ordering of a given partition ${\mathcal{P}}=\{P_{1},\ldots,P_{n}\}$ we can associate a function on $X$ defined by ${\mathcal{P}}(x)=j$ for $x\in P_{j}$.
@@
where the infimum is taken over all possible orderings for the partitions ${\mathcal{P}}$ and ${\mathcal{Q}}$. In case $|{\mathcal{P}}|\neq|{\mathcal{Q}}|$ we add empty cells to the partition with fewer elements in order to have two partitions with the same cardinality. If we identify partitions ${\mathcal{P}},{\mathcal{Q}}$ such that $d_{\mu}({\mathcal{P}},{\mathcal{Q}})=0$, then $d_{\mu}$ is indeed a metric on $\operatorname{\mathbf{P}}(X)$ (here we abuse the terminology and treat equivalence classes of partitions as partitions).
@@
In a previous work from 1981, Ratner [Rat81] used a definition of matching that requires the $(t,\varepsilon,{\mathcal{P}})$-matching to be an increasing function preserving the Lebesgue measure. Later, she proved that for ergodic flows the quantity $e(\Phi,u)$ remains the same if we use absolutely continuous or measure-preserving matching (Theorem 4 in [Rat17]), recovering a characterization of loosely Kronecker flows originally proved in 1981.
@@
The proof of Theorem 4.2 in [Rat81] relies on a characterization of loosely Kronecker flows due to Feldman [Fel80] which states that an ergodic flow is loosely Kronecker if and only if for any $\varepsilon>0$ and ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ we have $K_{t}(\varepsilon,{\mathcal{P}})=1$ (with respect to measure-preserving matchings) for every $t$ large enough.
Since the FK-pseudometric is a topological notion, we focus on partitions with good topological properties and we show that they are enough to prove $e(\Phi,u)=0$. In fact, we use the these partitions to show that for for any $\varepsilon>0$ and ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ we have $K_{t}(\varepsilon,{\mathcal{P}})=1$ (with respect to absolutely continuous matchings) for every $t$ large enough.
We start by fixing ${\mathcal{U}}$ a countable basis for the topology of $X$. We can assume without loss of generality that ${\mathcal{U}}$ is closed under finite unions and intersections. We say that a partition is ${\mathcal{U}}$-regular if at most one cell does not belong to ${\mathcal{U}}$.
@@
For a given $\mu\in{\mathcal{M}}(X)$, we say that a partition ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ is $\varepsilon$-essentially open if it is ${\mathcal{U}}$-regular and if there is a cell $P\in{\mathcal{P}}$ that does not belong to ${\mathcal{U}}$, then it satisfies $\mu(P)<\varepsilon$.
Besides having good topological properties, for each $\mu\in{\mathcal{M}}(X)$ the family of $\varepsilon$-essentially open partitions constitutes a dense set in the space of measurable partitions.
@@
Let ${\mathcal{P}}=\{P_{1},\ldots,P_{n}\}\in\operatorname{\mathbf{P}}(X)$. Fix $0<\delta<\varepsilon$. By regularity of $\mu$, for each $i\in\{1,\ldots,n\}$ there exists a compact $K_{i}\subseteq P_{i}$ such that $\mu(P_{i}\backslash K_{i})<\frac{\delta}{4n}$. Take $\{Q_{1},\ldots,Q_{n}\}\subseteq{\mathcal{U}}$ disjoint such that $K_{i}\subseteq Q_{i}$ and $\mu(Q_{i}\backslash K_{i})<\frac{\delta}{4n}$ for $i\in\{1,\ldots,n\}$. Set $Q_{n+1}=X\backslash\cup_{i=1}^{n}Q_{i}$ and ${\mathcal{Q}}=\{Q_{1},\ldots,Q_{n+1}\}$.
Observe that since $Q_{n+1}\subset\cup_{i=1}^{n}P_{i}\backslash K_{i}$ we have $\mu(Q_{n+1})<\frac{\delta}{4}<\varepsilon$ which means that ${\mathcal{Q}}$ is $\varepsilon$-essentially open. Moreover, $\mu(P_{j}\triangle Q_{j})<\frac{\delta}{2n}$ for every $j\in\{1,\ldots,n\}$ and therefore $d_{\mu}({\mathcal{P}},{\mathcal{Q}})<\delta$ which completes the proof. ∎
One direction of Theorem 1 is a direct consequence of García-Ramos and Kwietniak’s result (Theorem 4.5 in [GRK22]) together with some properties of the FK-pseudometric already presented.
@@
Assume that ${\mathcal{P}}$ and ${\mathcal{Q}}$ are ordered optimally with respect to $d_{\mu}$. Let $D=\{z\in X:{\mathcal{P}}(z)\neq{\mathcal{Q}}(z)\}$ and $C=X\backslash D$. For any $z\in X$ and $t\in\mathbb{R}^{+}$ we will denote by $C(z,t)=\{s\in[0,t]:\varphi^{s}(z)\in C\}$. Since $\mu$ is ergodic, applying the Birkhoff Ergodic Theorem to the characteristic function on $D$, we obtain that for $\mu$-almost every $z\in X$ it holds
[…]
So by Egorov’s Theorem there exists $H\subseteq\operatorname{Gen}(\mu)$ and $t_{0}>0$ such that $\mu(H)>1-\delta$ and for every $z\in H$ and $t\geq t_{0}$ we have
(4.1) $|\mu_{z,t}(D)-\mu(D)|=|\mu_{z,t}(D)-d_{\mu}({\mathcal{P}},{\mathcal{Q}})|<\frac{\delta}{2}.$
@@
(4.2) $\frac{\lambda(C(x,t))}{t}>1-d_{\mu}({\mathcal{P}},{\mathcal{Q}})-\frac{\delta}{2},$
and the same holds for $\lambda(C(y,t))$. Using these estimates together with the control of the measure by the derivative of $h$ (see condition (1) in Definition 4.1) and the fact that $\frac{\lambda(A_{0}^{\prime})}{t}>1-\varepsilon$ we have
(4.3) $\frac{\lambda(h^{-1}(A_{0}^{\prime}\cap C(y,t)))}{t}\geq\frac{\lambda(A_{0}^{\prime}\cap C(y,t))}{t}(1-\varepsilon)>1-2\varepsilon-d_{\mu}({\mathcal{P}},{\mathcal{Q}})-\frac{\delta}{2}.$
@@
The last ingredient is due to Ornstein, Rudolph and Weiss (see Chapter 7 in [ORW82]). The result was first obtained by Rudolph in [Rud78] for loosely Bernoulli flows (the positive entropy case) and a modification of the argument led to an analogous result for loosely Kronecker flows.
@@
Let $\Phi$ be a continuous flow on $X$ and $\mu\in\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$. The measure-preserving flow $(X,\Phi,\mu)$ is loosely Kronecker if and only if there exists a Borel set $H\subseteq X$ such that $\mu(H)=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
@@
By the discrete-time result analogous to Theorem 1 (Theorem 4.5 in [GRK22]) we obtain a Borel set $H\subseteq X$ with $\mu(H)=1$ such that $\operatorname{\rho_{FK}}(x,y)=0$ with respect to $\varphi^{t}$ for every $x,y\in H$ By Proposition 3.17 we obtain $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
For the converse we fix $\varepsilon>0$, a partition ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ and a Borel set $H_{0}\subseteq X$ such that $\mu(H_{0})=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H_{0}$.
From the density of essentially open partitions there exists an $\varepsilon$-essentially open partition ${\mathcal{Q}}=\{Q_{0},\ldots,Q_{n}\}$ such that $d_{\mu}({\mathcal{P}},{\mathcal{Q}})<\varepsilon$. We assume without loss of generality that $Q_{0}$ is the compact cell and take an open set $V_{0}$ containing $Q_{0}$ such that $\mu(V_{0})<\varepsilon$.
Let $0<\delta<\varepsilon$ be a Lebesgue number for the open cover ${\mathcal{W}}=\{V_{0},Q_{1},\ldots,Q_{n}\}$ of $X$.
Since $\mu$ is ergodic, there exist a compact set $H_{1}\subseteq\operatorname{Gen}(\mu)\cap H_{0}$ and $t_{1}>1$ such that $\mu(H_{1})>1-\delta$ and $\mu_{x,t}(V_{0})<\varepsilon$ for every $t\geq t_{1}$ and $x\in H_{1}$.
Fix $x_{0}\in H_{1}$. We observe that $\tilde{f}_{t,\delta}(x_{0},\cdot)$ is continuous for every $t\geq 1$, so $\tilde{f}_{\delta}(x_{0},\cdot)$ is measurable. Thus there exist a compact set $H_{2}\subseteq H_{1}$ and $t_{2}\geq t_{1}$ such that $\mu(H_{2})>1-\delta$ and $\tilde{f}_{t,\delta}(x_{0},y)<\delta$ for every $t\geq t_{2}$ and $y\in H_{2}$.
For $t\geq t_{2}$ and $y\in H_{2}$ let $h\colon A_{0}\to A_{0}^{\prime}$ be a $\widetilde{(t,\delta,\delta)}$-matching between $x_{0}$ and $y$. Set $A\subseteq A_{0}$ as
@@
Using an argument similar to the one in the proof of Lemma 4.5 together with the fact that $\delta<\varepsilon$ we can see that $h|_{A}\colon A\to h(A)$ is a $(t,3\varepsilon,{\mathcal{Q}})$-matching between $x_{0}$ and $y$.
Lastly, we apply Lemma 4.5 and assume that the obtained set $H$ is contained in $H_{2}$ and $t_{0}\geq t_{2}$. So, for every $y\in H$ and $t\geq t_{0}$ we have
@@
This means that $H\subseteq B_{t}(x_{0},15\varepsilon,{\mathcal{P}})$, and consequently, $K_{t}(15\varepsilon,{\mathcal{P}})=1$ for every $t\geq t_{0}$. Since ${\mathcal{P}}\in\operatorname{\mathbf{P}}(X)$ and $\varepsilon>0$ were arbitrary, we have $e(\Phi,u)=0$ for every $u\in U$. Using Theorem 4.2 we conclude that $(X,\Phi,\mu)$ is loosely Kronecker. ∎